OpenStudy (anonymous):

Will Fan And Medal What could you do to eliminate the fractions in the equation c-sevenths minus three-sevenths equals one-seventh? A. Add three-sevenths to both sides of the equation. B. Multiply both sides of the equation by one-seventh. C. Multiply both sides of the equation by 7. D. Subtract three-sevenths from both sides of the equation.

OpenStudy (anonymous):

c/7 - 3/7 = 1/7 To eliminate the fraction, you must multiply the entire equation by 7 (7) c/7 - (7)3/7 = (7)1/7 c - 3 = 1 c = 1 + 3 c = 4 ===================== 5 4/9 - (-2 8/9) -- negative times a negative = a positive, thus changing your problem to : 5 4/9 + 2 8/9 -- change to improper fractions 49/9 + 26/9 = 75/9 turned into a mixed number = 8 1/3